Discount Amortization

On the first day of the fiscal year, a company issues a $3,000,000, 11%, 9-year bond that pays semiannual interest of $165,000 ($3,000,000 × 11% × ½), receiving cash of $2,547,343.

Journalize the first interest payment and the amortization of the related bond discount. Round to the nearest dollar. If an amount box does not require an entry, leave it blank.

Interest Expense 210265 (incorrect)
Discount on Bonds Payable 45265 (incorrect)
Cash 165000 (correct)

I tried rounding One number up for Interest expense and Discount Payable up but still is wrong.

Premium Amortization

On the first day of the fiscal year, a company issues a $2,900,000, 10%, 6-year bond that pays semiannual interest of $145,000 ($2,900,000 × 10% × ½), receiving cash of $3,320,357.

Journalize the first interest payment and the amortization of the related bond premium. Round to the nearest dollar. If an amount box does not require an entry, leave it blank.

Interest Expense
Premium on Bonds Payable   
Cash 145000 (correct)
Answer #1

Part 1.

Discount amortization :

Par value of bonds = $3,000,000

Cash proceeds from issue of bonds = $2,547,343

Discount on bonds payable = Par value of bonds - Cash proceeds from issue of bonds

= 3,000,000 - 2,547,343

= $452,657

Since it is a 9 year bond and interest is payable semi - annually, hence bond discount will be amortized in 18 semi annual installments.

Semi annual amortization of bond discount = 452,657/18

= $25,148

Journal

Account Title and Explanation

Debit

Credit

Interest expense 190,148
Discount on Bonds Payable 25,148
Cash 165,000

Part 2.

premium amortization :

Par value of bonds = $2,900,000

Cash proceeds from issue of bonds = $3,320,357

Premium on bonds payable = Cash proceeds from issue of bonds - Par value of bonds

= 3,320,357 - 2,900,000

= $420,357

Since it is a 6 year bond and interest is payable semi - annually, hence bond discount will be amortized in 12 semi annual installments.

Semi annual amortization of bond discount = 420,357/12

= $35,030

Journal

Account Title and Explanation

Debit

Credit

Interest expense 109,970
Premium on Bonds Payable 35,030
Cash 145,000

kindly give a positive rating if you are satisfied with the solution. do comment if you have any query, Thanks.